OpenStudy (anonymous):

Fun question If f(0) = f(1) = 0 and f" exists, show that

OpenStudy (anonymous):

\[\int\limits_{0}^{1} f''(x)f(x)dx = - \int\limits_{0}^{1} [f'(x)]^2 dx\] shouldn't be too bad :)

OpenStudy (anonymous):

LOL

OpenStudy (jhannybean):

Alright what if you combine \(f''(x)f(x) \implies (f'''(x)dx)\) then do the chain rule twice?

OpenStudy (jhannybean):

something like that.

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

My initial thought:\[ (ff')' = f'f'+ff'' \]

OpenStudy (jhannybean):

That's where I was going with it.

OpenStudy (jhannybean):

Guess triple prime wasn't needed.

ganeshie8 (ganeshie8):

by parts !

OpenStudy (anonymous):

wio U are right

OpenStudy (anonymous):

(ff')=(f')^2+ff"

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{1} f''(x)f(x)dx = f'(x)f(x)\Bigg|_{0}^1 - \int \limits_0^1 f'(x) f'(x) dx\]

OpenStudy (anonymous):

We know that: \[ \int_0^2 (f(x)f'(x))'dx=\int_0^1d\bigg(f(x)f'(x)\bigg)= 0f'(2)-0f'(0)= 0 \]

OpenStudy (anonymous):

And oo:\[\int\limits_{0}^{1} f''(x)f(x)dx +\int\limits_{0}^{1} [f'(x)]^2 dx = 0\]

OpenStudy (anonymous):

and so

OpenStudy (anonymous):

@iambatman Is this enough?

OpenStudy (anonymous):

Yes, that's enough haha, maybe next time I'll have a no ganeshie rule! :P

OpenStudy (anonymous):

What? I came up with the anti-derivative first thought...

OpenStudy (anonymous):

though Laptop keyboard is unreliable

ganeshie8 (ganeshie8):

lol ask tough questions instead

OpenStudy (bibby):

shots fired

OpenStudy (anonymous):

It wasn't suppose to be tough :(, just fun

OpenStudy (jhannybean):

pew pew

OpenStudy (anonymous):

Guess, I'll find some questions related to eliassaab's haha, if you want a real challenge :P

ganeshie8 (ganeshie8):

I think we exceeded the quota for number theory this week, im eager to do more calc post more questions batman xD

OpenStudy (anonymous):

Haha, well mhm, I'll have to find/ think of some tricky ones for calc...I will see :D, and I agree there have been many number theory questions the past week D:

OpenStudy (mathmath333):

does\( f(x)=xsin(\pi x)\) satisfies it

ganeshie8 (ganeshie8):

is it twice differentiable ?

OpenStudy (mathmath333):

yes i think

ganeshie8 (ganeshie8):

yes it satisfies the hypothesis so both integrals need to evaluate to same numbers

OpenStudy (anonymous):

That's neat haha

OpenStudy (jhannybean):

Post questions involving Vieta's Formula so I can LEARN how to use it!!!!

ganeshie8 (ganeshie8):

was yesterday's drill useful jhanny ?

OpenStudy (jhannybean):

....Hah. <_<

OpenStudy (anonymous):

Haven't heard of that in a while lol